OpenStudy (secret-ninja):

Medals. Which of the following best defines 2 to the power of 2 over 3? Cube root of 4 Square root of 4 Cube root of 2 Square root of 2

OpenStudy (secret-ninja):

\[2\frac{ 2 }{ 3 }\]

OpenStudy (secret-ninja):

@ParthKohli @zepdrix Can you guys help?

OpenStudy (solomonzelman):

\(\large\color{black}{ 2^{2/3} }\) There is a rule, \(\large\color{black}{ A^{B/C}=\sqrt[C]{A^{B~}} }\) Use this rule to re-write \(\large\color{black}{ 2^{2/3} }\).

OpenStudy (anonymous):

i think it's A

OpenStudy (secret-ninja):

\[\sqrt[3]{2}^2\]

OpenStudy (secret-ninja):

Like that?

OpenStudy (anonymous):

no

OpenStudy (solomonzelman):

yes

OpenStudy (anonymous):

\[\sqrt[3]{2^2?}\]

OpenStudy (solomonzelman):

it is the same thing

OpenStudy (solomonzelman):

whether 2 is inside the root or outside the root.

OpenStudy (solomonzelman):

you can write it like this, and +, she meant the 2 to be inside the root most likely.

OpenStudy (anonymous):

so the first one is correct

OpenStudy (secret-ninja):

I did, i'm just not used to using it lol sorry >.<

OpenStudy (solomonzelman):

So when you have \(\large\color{black}{ \sqrt[3]{\frak number} }\) it is a cube root of a number.

OpenStudy (solomonzelman):

and when you re-write \(\large\color{black}{ 2^2 }\) that you got inside, you get 4. So it would be just \(\large\color{black}{ \sqrt[3]{4} }\) .

OpenStudy (solomonzelman):

Which option do you pick ?

OpenStudy (secret-ninja):

So the answer is A?

OpenStudy (anonymous):

yes

OpenStudy (solomonzelman):

yes, there we go:)

OpenStudy (secret-ninja):

Thanks!. :)

OpenStudy (solomonzelman):

Anytime:)